40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 13 23:58:38 2011, Koji, Update, LSC, fixing REFL165  PA131612.jpgPA131618.jpg
    Reply  Sat Oct 15 16:42:08 2011, Koji, Update, LSC, Testing REFL165  REFL165_test_111014_KA.pdfREFL165_transimpedance2.pdf
       Reply  Sat Oct 15 17:06:20 2011, Koji, Update, LSC, Installation REFL165  
       Reply  Sun Oct 16 02:30:00 2011, rana, Update, Electronics, Testing REFL165  
    Reply  Wed Mar 1 17:25:28 2017, Steve, Update, LSC, MCREFL condition pictures IMG_3646.JPGmcRefl_1.jpgmcRefl_3.jpgmcRefl_5.jpg
       Reply  Fri Mar 17 14:49:09 2017, gautam, Update, LSC, MCREFL condition pictures IMG_7219.JPGMCREFL_X.pdfMCREFL_Y.pdf
Message ID: 5672     Entry time: Sat Oct 15 17:06:20 2011     In reply to: 5671
Author: Koji 
Type: Update 
Category: LSC 
Subject: Installation REFL165  

REFL165 was installed on the AP table last night.

Meanwhile I found the DC power level at the REFL PDs were 0.8~1.2V if the PRM is aligned and the IFO is not locked.
This corresponds to 16~24mA (20~30mW). This is too big.

The HWP of the REFL path were adjusted so that we have 6~10mA (8~12mW) on each PDs.

ELOG V3.1.3-